# FIGRET: Fine-Grained Robustness-Enhanced Traffic Engineering

### v1

@@ 4. FIGRET design
In this section, we present the design of FIGRET. In §4.1, we introduce how FIGRET addresses traffic bursts while solving TE solutions based on historical traffic. Subsequently, §4.2 discusses the transformation of problem-solving into the design of loss functions using deep learning. The design of the loss function and the network architecture are then detailed in §4.3 and §4.4, respectively. Finally, §4.5 also covers how FIGRET handles link failures.
@@ 4.1. Design for managing traffic bursts
By defining a set to account for traffic bursts. A prevalent method involves defining the mismatch as a set $\Delta$. During the optimization process, the sum of a specific mismatch $\delta$ from $\Delta$ and $D^{\text{expect}}$ is used as an estimate for the actual traffic, denoted as $\hat{D_{t}}=\delta+D^{\text{expect}}$. TE configurations are then determined based on $\hat{D_{t}}$, i.e.,
@@
Drawback: Adjusting the size of the set $\Delta$ allows for the modulation of the algorithm’s emphasis on traffic bursts. However, defining this set accurately is quite challenging. A larger set size implies that TE takes into account a greater traffic variability, which, while potentially increasing the algorithm’s robustness, may diminish its focus on $D^{\text{expect}}$. Conversely, a smaller set size means fewer traffic variable scenarios are considered, potentially compromising robustness. Therefore, this approach is not considered in our design.
By imposing constraints to account for traffic bursts. When considering both $D^{\text{expect}}$ and $\Delta$, the utilization of each edge can be expressed as: $\sum_{s,d\in V,p\in P_{sd},e\in p}(D_{sd}^{\text{expect}}+\delta_{sd})\cdot r_{p}/c(e)$. Consequently, the actual impact of $\delta_{sd}$ on each edge can be represented by its coefficient. To mitigate this impact, it is essentially about ensuring that this coefficient is less than a certain constraint. The degree of robustness can be controlled by adjusting the stringency of this constraint.
As discussed in §2, we aim to impose varying degrees of robustness on different paths based on the traffic characteristics of the SD pairs. Therefore, this constraint should be a function related to the SD pairs, defined as $\mathcal{F}:(s,d)\rightarrow\mathbb{R}^{+}$.
(3) $\forall s,d\in V,\forall p\in P_{sd},\forall e\in P\quad\frac{r_{p}}{c(e)}\leq\mathcal{F}((s,d))$
Considering the definitions of path capacity $C_{p}$, as expressed by $C_{p}=\min_{e\in p}c(e)$, it follows that Equation 3 is equivalent to Equation 4.
(4) $\forall{s,d\in V},\forall{p\in P_{st}}\quad\frac{r_{p}}{C_{p}}\leq\mathcal{F}((s,d))$
As indicated in Equation 4, ensuring that the ratio $\frac{r_{p}}{C_{p}}$ falls within certain constraints guarantees the robustness of network configurations against burst traffic. Therefore $\frac{r_{p}}{C_{p}}$ can be used as a metric to represent the degree of robustness in the network configuration. So we define path sensitivity $\mathscr{S}_{p}$ for path $p$ as
$\mathscr{S}_{p}=\frac{r_{p}}{C_{p}}$
By taking into account the metric of path sensitivity when specifying TE configurations and ensuring it meet certain constraints, TE can be aided in minimizing the impact of $D^{\text{burst}}$ on network performance. Therefore, our solution can be expressed as Equation 5.
@@
Remark. In Equation 5, both $D_{t}^{\text{expect}}$ and $\mathcal{F}$ are determined based on historical traffic data $\{D_{1},\dots,D_{t-1}\}$. Selecting an appropriate $D_{t}^{\text{expect}}$ and $\mathcal{F}$ for a specific network topology and traffic pattern is not trivial. Existing methods have not finely discriminated between different SD pairs. For example:
  • Traditional methods based on direct optimization of MLU after prediction, as well as those based on deep learning (Valadarsky et al., 2017; Perry et al., 2023), set $\mathcal{F}((s,t))\equiv+\infty,\forall s,t\in V$. They assume that the real flow can be entirely represented by $D^{\text{expect}}$. However, due to the inherent dynamic nature of traffic demands, mismatches are inevitable.
  • Desensitization-based TE schemes (Poutievski et al., 2022) set $\mathcal{F}((s,d))\equiv\text{const},\forall s,d\in V$. A single constant value makes it challenging to perfectly balance burst resistance and average performance. Another approach (Teh et al., 2022) sets $\mathcal{F}((s,d))$ as an objective, aiming to minimize the maximum $\mathcal{F}((s,d))$ among all SD pairs. However, it also does not account for differences among SD pairs. Even if a stable SD pair has a larger $\mathcal{F}((s,d))$, it would not impact robustness.
@@ 4.2. FIGRET framework
In this section, we explain the overall framework of FIGRET for solving Equation 5. To solve Equation 5, we first consider the following: both $D_{t}^{\text{expect}}$ and $\mathcal{F}$ are determined based on historical traffic data $\{D_{1},\dots,D_{t-1}\}$. If the mapping relationship between $\{D_{1},\dots,D_{t-1}\}$ and both $D_{t}^{\text{expect}}$ and $\mathcal{F}$ is already known, then Equation 5 can be directly solved. Therefore, the framework for solving Equation 5 is divided into two types based on whether an explicit $D_{t}^{\text{expect}}$ and $\mathcal{F}$ are pre-determined.
The first approach involves, initially, attempting to explicitly solve for $D_{t}^{\text{expect}}$ and $\mathcal{F}$, followed by employing a linear programming-based algorithm to solve the Equation 5. However, this method encounters two main issues. First, there is a mismatch (Perry et al., 2023) between predicting $D_{t}^{\text{expect}}$ accurately and the ultimate goal of TE, which is to minimize MLU. This mismatch arises due to the influence of network topology and edge capacities; accurately predicting certain flows is more crucial than others. For instance, accurately predicting traffic for SD pairs with paths of high capacity is less critical. But the objective of accurate prediction (usually employing mean-squared-error, etc., as the loss) does not account for this. An example is illustrated in Figure 16. As for $\mathcal{F}$, it is even more challenging to define a specific approach to find an optimal $\mathcal{F}$. Second, using linear programming introduces substantial computational delays, and the algorithm’s scalability to network topology is not guaranteed. Therefore, the approach of first solving for $D_{t}^{\text{expect}}$ and $\mathcal{F}$ before solving the problem is not ideal.
An alternative approach eschews the explicit computation of $D_{t}^{\text{expect}}$ and $\mathcal{F}$. Instead, it directly establishes an end-to-end relationship between historical traffic data $\{D_{1},\dots,D_{t-1}\}$ and a TE configuration $R_{t}$ that aims to minimize MLU and assures robustness. The fundamental effectiveness of this method lies in its intrinsic capability to implicitly weigh the importance of different SD pairs based on network topology and capacity. This effectively mitigates the previously observed mismatch between obtaining $D_{t}^{\text{expect}}$ and $\mathcal{F}$, and the ultimate goal of minimizing MLU. Based on the above considerations, we also adopt this approach in FIGRET.
Leveraging Deep Learning Although we recognize the need to establish an end-to-end relationship between historical traffic data $\{D_{1},\ldots,D_{t-1}\}$ and a TE configuration $R_{t}$ that minimizes MLU and assures robustness, the complexity of network topologies and real-world traffic patterns defy straightforward mathematical characterization. In response, FIGRET adopts a deep neural network (DNN) approach, effectively sidestepping the complexities of directly solving Equation 5 by reframing the challenge into a well-designed loss function that captures the essence of the optimization problem. The powerful learning capabilities of DNNs are leveraged to effectively approximate the solution to Equation 5, accommodating the dynamic and uncertain nature of network traffic and topology.
@@ 4.3. Design loss function of FIGRET
In FIGRET, a Deep Neural Network (DNN) is employed to establish a mapping between historical traffic data, $\{D_{1},\dots,D_{t-1}\}$, and TE configurations. The input to FIGRET comprises historical traffic data, while its output is a TE configuration. To maintain a consistent input size for the DNN and prevent an overload of traffic data when $t$ is large, a temporal window $H$ is typically chosen (Valadarsky et al., 2017; Perry et al., 2023). This leads to the use of traffic data within this window, $\{D_{t-H},\dots,D_{t-1}\}$, as the input. During training, FIGRET receives $\{D_{t-H},\dots,D_{t-1}\}$, and subsequently outputs a TE Configuration $R_{t}$. We denote this output as $R_{t}=\pi_{\theta}(D_{t-H},\dots,D_{t-1})$, where $\pi$ represents the mapping function from DNN inputs to outputs, and $\theta$ denotes the DNN’s link weights. The revealed $D_{t}$ allows for the calculation of the loss for this $R_{t}$ through the loss function $\mathcal{L}(R_{t},D_{t})$, after which a gradient descent optimizer updates the parameters $\theta$. Thus, to enable FIGRET to approximate the solution to Equation 5, a well-designed loss function is essential. This function must capture the essence of the optimization problem, guiding the neural network towards effective and efficient solutions.
Loss function design. The design of our loss function is intrinsically aligned with the Lagrangian relaxation of Equation 5, as delineated in Equation 6.
@@
This relaxation framework guides us in structuring a loss function that encompasses two components. The first component corresponds to minimizing max link utilization.
(7) $\mathcal{L}_{1}=M(R_{t},D_{t}),$
@@
With the first loss function in place, upon completion of training, FIGRET implicitly learns a probability distribution $P(D_{t}^{\text{expect}}|D_{t-1},\dots,D_{t-H})$. It can then output an $R_{t}$ that optimizes $\mathbb{E}_{D_{t}^{\text{expect}}}[M(\mathcal{R}^{t},D_{t}^{\text{expect}})]$ with respect to $P(D_{t}^{\text{expect}}|D_{t-1},\dots,D_{t-H})$ ($D_{t}^{\text{expect}}$ can be multiple). This illustrates that optimizing MLU directly, rather than focusing on the accuracy of $P(D_{t}|D_{t-1},\dots,D_{t-H})$, essentially employs network topology and capacity information to weight $P(D_{t}|D_{t-1},\dots,D_{t-H})$. This approach effectively resolves the mismatch between prediction accuracy and the objective of minimizing MLU.
@@
The second component of the loss function reflects the constraints imposed by path sensitivities. We use $\sigma^{2}_{D_{sd},[1-T]}$ to denote the variance of traffic demands $D_{sd}$ from source $s$ to destination $d$ within the time range from 1 to $T$. We denote $\mathscr{S}_{sd}^{\text{max}}$ as the maximum path sensitivity among all paths serving from source $s$ to destination $d$ in the configuration $\mathcal{R}_{t}$. Based on this, we can express $\mathcal{L}_{2}$ as Equation 8.
(8) $\mathcal{L}_{2}=\sum_{\forall s,d\in V}\sigma^{2}_{D_{sd},[1-T]}\times\mathscr{S}_{sd}^{\text{max}},$
@@
The rationale behind the formulation of $\mathcal{L}_{2}$ is rooted in the differential treatment of path sensitivities across various source-destination (SD) pairs, guided by the constraints $\mathcal{F}((s,d))$. In essence, for SD pairs exhibiting stable traffic patterns, denoted as $(s,d)_{\text{stable}}$, the model enforces less stringent path sensitivity constraints. Conversely, for SD pairs characterized by bursty traffic dynamics, labeled as $(s,d)_{\text{bursty}}$, the model imposes stricter path sensitivity restrictions. This distinction is encapsulated in the relationship $\mathcal{F}((s,d)_{\text{stable}})>\mathcal{F}((s,d)_{\text{bursty}})$. In other words, assuming in a given scenario $\mathscr{S}_{sd_{\text{bursty}}}^{\text{max}}=\mathscr{S}_{sd_{\text{stable}}}^{\text{max}}$, the bursty traffic SD pair might incur a greater penalty. To ensure $\mathcal{L}_{2}$ is designed to meet this criterion, we calculate the variance of traffic demand $\sigma^{2}_{D_{sd},[1-T]}$ for each SD pair from time 1 to $T$, representing the degree of variation in the flow. This provides a quantitative measure of traffic fluctuation for the network. Subsequently, we weight these variances against the maximum path sensitivity $\mathscr{S}_{sd}^{\text{max}}$ of paths serving the corresponding SD pair. This approach imposes stricter sensitivity penalties on flows with higher variance and lighter penalties on those with lower variance. Such a weighted method effectively differentiates between various traffic characteristics, making corresponding adjustments to the impact on network performance.
@@
When training the neural network, we consider both the loss term MLU $\mathcal{L}_{1}$ and the plenty term $\mathcal{L}_{2}$. This comprehensive consideration aims to guide the network in learning how to adjust path sensitivity constraints for different SD pairs while maintaining MLU optimization. By this method, we ensure that the network not only meets the MLU optimization objectives but also enhances the burst resistance of different SD pairs in a fine-grained manner.

### v3

@@ 4. FIGRET design
In this section, we present the design of FIGRET. In §4.1, we introduce how FIGRET addresses traffic bursts while solving TE solutions based on historical traffic. Subsequently, §4.2 discusses the transformation of problem-solving into the design of loss functions using deep learning. The design of the loss function and the deep neural network architecture are then detailed in §4.3 and §4.4, respectively. Finally, §4.5 discusses how FIGRET handles node/link failures.
@@ 4.1. Design for managing traffic bursts
By defining a set to account for traffic bursts. A prevalent method involves defining the bursts as a set $\Delta$. During the optimization process, the sum of a specific burst $\delta\in\Delta$ and $D^{\text{expect}}$ is used as an estimate for the actual traffic, denoted as $\hat{D_{t}}=\delta+D^{\text{expect}}$. The objective is to find a TE configuration that offers the best worst-case guarantee for all the $\delta\in\Delta$, i.e.,
@@
Drawback: The above robust optimization based approach could offer varying levels of robustness with different choices of $\Delta$, but still faces the following shortcomings:
  • It is hard to find the most appropriate $\Delta$ that yields the best trade-off between normal case performance and burst case performance. Although one could use Chebyshev’s inequality (Marshall and Olkin, 1960) to quantify a $\Delta$ such that the probability $P(D_{t}\notin\hat{\mathcal{D}})$ is no larger than a small value $\alpha$, $D_{t}$ may still go beyond $\hat{\mathcal{D}}$ and the performance guarantee offered by Equation (2) becomes invalid. Admittedly, we could increase $\Delta$ such that the probability $\alpha$ is extremely small, but the normal-case performance could then be compromised.
  • The objective function in Equation (2) aims to optimize the worst case within the set $\hat{\mathcal{D}}$. Note that $\hat{\mathcal{D}}$ may contain an infinite number of distinct traffic patterns, dramatically increasing the computational complexity of solving Equation (2). In contrast, our chosen anti-burst strategy in Equation (5) only requires optimization for a single traffic pattern $D_{t}^{\text{expect}}$.
By imposing constraints to account for traffic bursts. When considering both $D^{\text{expect}}$ and $\delta$, the utilization of each edge can be expressed as: $\sum_{s,d\in V,p\in P_{sd},e\in p}(D_{sd}^{\text{expect}}+\delta_{sd})\cdot r_{p}/c(e)$. Consequently, the actual impact of $\delta_{sd}$ on each edge can be represented by its coefficient. To mitigate the impact of traffic bursts, we could impose an upper bound on this coefficient. The degree of robustness can be controlled by adjusting the upper bound of each coefficient.
As discussed in §2, we aim to impose varying levels of robustness on different paths based on the traffic characteristics of the SD pairs. Therefore, this constraint should be a function related to the SD pairs, defined as $\mathcal{F}:(s,d)\rightarrow\mathbb{R}^{+}$.
(3) $\forall s,d\in V,\forall p\in P_{sd},\forall e\in P\quad\frac{r_{p}}{c(e)}\leq\mathcal{F}((s,d)).$
Considering the definitions of path capacity $C_{p}$, as expressed by $C_{p}=\min_{e\in p}c(e)$, it follows that Equation (3) is equivalent to Equation (4).
(4) $\forall{s,d\in V},\forall{p\in P_{st}}\quad\frac{r_{p}}{C_{p}}\leq\mathcal{F}((s,d)).$
As indicated in Equation (4), ensuring that the ratio $\frac{r_{p}}{C_{p}}$ falls below a certain limit guarantees the robustness of network configurations against burst traffic. Therefore, $\frac{r_{p}}{C_{p}}$ can be used as a metric to represent the level of robustness in the network configuration. So we define path sensitivity $\mathscr{S}_{p}$ for path $p$ as
$\mathscr{S}_{p}=\frac{r_{p}}{C_{p}}.$
By taking into account the path sensitivity metric in our TE formulation, the impact of traffic bursts can be mitigated. The resulting formulation is shown below:
@@
Remark. In Equation (5), both $D_{t}^{\text{expect}}$ and $\mathcal{F}$ are determined based on historical traffic data $\{D_{1},\dots,D_{t-1}\}$. Selecting an appropriate $D_{t}^{\text{expect}}$ and $\mathcal{F}$ for a specific network topology and traffic pattern is not trivial. Existing methods have not finely discriminated between different SD pairs. For example:
  • Traditional methods based on direct optimization of MLU after prediction, as well as those based on deep learning (Valadarsky et al., 2017; Perry et al., 2023), set $\mathcal{F}((s,t))\equiv+\infty,\forall s,t\in V$. They assume that the real flow can be entirely represented by $D^{\text{expect}}_{t}$. However, due to the inherent dynamic nature of traffic demands, mismatches are inevitable.
  • Desensitization-based TE schemes (Poutievski et al., 2022) set $\mathcal{F}((s,d))\equiv\text{const},\forall s,d\in V$. A single constant value makes it challenging to perfectly balance burst resistance and average performance. Another approach (Teh et al., 2022) sets $\mathscr{S}_{p}$ as an objective, aiming to minimize the maximum $\mathscr{S}_{p}$ among all paths. However, it also does not account for the difference among different SD pairs.
@@ 4.2. FIGRET framework
4.2.1. How to compute a TE solution
Given the FIGRET formulation in Equation (5), the next challenge is to efficiently compute TE solutions. Two approaches to consider are the two-stage method and the end-to-end method.
The two-stage method first explicitly estimates $D_{t}^{\text{expect}}$ and $\mathcal{F}$. Then, it employs a linear programming-based algorithm to solve Equation (5). However, this approach has the following shortcomings, making it far from ideal:
  • Due to the existence of highly bursty source-destination pairs, predicting a suitable $D_{t}^{\text{expect}}$ is challenging.
  • There exists a mismatch between the upstream tasks of determining $D_{t}^{\text{expect}}$ and $\mathcal{F}$, and the downstream task of optimizing MLU. Predicting $D_{t}^{\text{expect}}$ often relies on Mean Squared Error (MSE) for evaluation, which does not account for network topology—an essential factor in MLU optimization. For instance, in SD pairs with high-capacity paths, the precision of traffic predictions becomes less critical, as illustrated in Figure LABEL:fig:_mismatch_in_prediction in Appendix G.1. This oversight can adversely affect TE performance. A similar mismatch affects $\mathcal{F}$, where constraints should reflect not only traffic characteristics but also the topological structure.
  • Employing linear programming involves high computational complexity and may not scale to large networks.
The end-to-end method avoids the explicit prediction of $D_{t}^{\text{expect}}$ and $\mathcal{F}$. Instead, it directly establishes a relationship between historical traffic data $\{D_{1},\dots,D_{t-1}\}$ and a TE configuration $R_{t}$, with a goal to minimize MLU and ensure robustness. This method’s advantage lies in its ability to consider the impact of network topology, eliminating the mismatch between the upstream traffic prediction and the downstream network optimization (Perry et al., 2023). Thus, its performance is better than that of the two-stage method (See Appendix C for more details).
Overall, based on the above considerations, we opt for the end-to-end method to compute TE solutions.
4.2.2. Leveraging deep learning to implement the end-to-end method
We employ an end-to-end method to establish a mapping between historical traffic data $\{D_{1},\ldots,D_{t-1}\}$ and TE configuration $R_{t}$ that minimizes MLU while ensuring robustness. Since this end-to-end method does not pre-solve $D_{t}^{\text{expect}}$, we are unable to establish a clear optimization problem as in Equation (5). Therefore, we cannot utilize a linear programming-based approach to implement the end-to-end method. In response, drawing inspiration from DOTE (Perry et al., 2023), FIGRET adopts a deep neural network (DNN) to implement the end-to-end method. By designing a well-designed burst-aware loss function, FIGRET can output TE solutions that effectively balance normal and burst traffic scenarios.
@@ 4.3. Design loss function of FIGRET
In FIGRET, a Deep Neural Network (DNN) is employed to establish a mapping between historical traffic data $\{D_{1},\dots,D_{t-1}\}$ and TE configurations. To maintain a consistent input size for the DNN and prevent an overload of traffic data when $t$ is large, a temporal window $H$ is typically chosen (Valadarsky et al., 2017; Perry et al., 2023). During training, FIGRET receives $\{D_{t-H},\dots,D_{t-1}\}$, and subsequently outputs a TE Configuration $R_{t}$. We denote this output as $R_{t}=\pi_{\theta}(D_{t-H},\dots,D_{t-1})$, where $\pi$ represents the mapping function from DNN inputs to outputs, and $\theta$ denotes the DNN’s link weights. The revealed $D_{t}$ allows for the calculation of the loss for this $R_{t}$ through the loss function $\mathcal{L}(R_{t},D_{t})$, after which a gradient descent optimizer updates the parameters $\theta$. Thus, to enable FIGRET to output TE solutions with fine-grained robustness, a well-designed loss function is essential.
The design of our loss function is intrinsically aligned with the Lagrangian relaxation of Equation (5), as delineated in Equation (6). This relaxation framework guides us in structuring a loss function that encompasses two components.
@@
4.3.1. Loss for MLU
The first component corresponds to minimizing max link utilization.
(7) $\mathcal{L}_{1}=M(R_{t},D_{t}).$
@@ 4.3.1. Loss for MLU
With the first loss function in place, upon completion of training, FIGRET implicitly learns a probability distribution $P(D_{t}^{\text{expect}}|D_{t-1},\dots,D_{t-H})$. It can then output an $R_{t}$ that optimizes $\mathbb{E}_{D_{t}^{\text{expect}}}[M(\mathcal{R}^{t},D_{t}^{\text{expect}})]$ with respect to $P(D_{t}^{\text{expect}}|D_{t-1},\dots,D_{t-H})$.
4.3.2. Loss for fine-grained robustness
The second component of the loss function reflects the constraints imposed by path sensitivities. Since we adopt an end-to-end TE approach without explicitly solving for $\mathcal{F}((s,d))$ beforehand, we can not directly use the second term of Equation (6)—$\sum_{s,d\in V}\sum_{p\in P_{sd}}\lambda_{sd}(\mathscr{S}_{p}-\mathcal{F}((s,d)))$—as the loss function. In response, we design the following heuristic scheme for constructing the second component of the loss.
@@ 4.3.2. Loss for fine-grained robustness
How to calculate $\mathcal{L}_{2}$? We use $\sigma^{2}_{D_{sd},[1-T]}$ to denote the variance of traffic demands $D_{sd}$ from source $s$ to destination $d$ over the training period from time 1 to $T$. We denote $\mathscr{S}_{sd}^{\text{max}}$ as the maximum path sensitivity among all paths serving from source $s$ to destination $d$ in the configuration $\mathcal{R}_{t}$. Based on this, we can express $\mathcal{L}_{2}$ as Equation (8).
(8) $\mathcal{L}_{2}=\sum_{\forall s,d\in V}\sigma^{2}_{D_{sd},[1-T]}\times\mathscr{S}_{sd}^{\text{max}}.$
@@
How does Equation (8) capture fine-grained robustness? The original term $\sum_{s,d\in V}\sum_{p\in P_{sd}}\lambda_{sd}(\mathscr{S}p-\mathcal{F}((s,d)))$ facilitates fine-grained robustness because the values of $\mathcal{F}((s,d))$ differ among various source-destination pairs, resulting in different penalties for the same path sensitivity across different paths. To elaborate, for SD pairs with stable traffic patterns, denoted as $(s,d)_{\text{stable}}$, and those characterized by bursty traffic dynamics, labeled as $(s,d)_{\text{bursty}}$, the constraints are more stringent for bursty SD pairs, i.e., $\mathcal{F}((s,d)_{\text{stable}})>\mathcal{F}((s,d)_{\text{bursty}})$. Therefore, according to $\mathscr{S}p-\mathcal{F}((s,d)))$, the path serving for bursty traffic SD pair incurs a greater penalty. Our design of $\mathcal{L}_{2}$ achieves a similar functionality. We calculate the variance of traffic demand $\sigma^{2}_{D_{sd},[1-T]}$ for each SD pair from time 1 to $T$, representing the degree of variation in the flow. This provides a quantitative measure of traffic fluctuation for the network. Subsequently, we weight these variances with the maximum path sensitivity $\mathscr{S}_{sd}^{\text{max}}$ of paths serving the corresponding SD pair. This approach imposes stricter sensitivity penalties on flows with higher variance and lighter penalties on those with lower variance.
@@
When training the neural network, we consider both the loss term MLU $\mathcal{L}_{1}$ and the robustness enhancement term $\mathcal{L}_{2}$. This comprehensive consideration aims to guide the network in learning how to adjust path sensitivity constraints for different SD pairs while maintaining MLU optimization. By this method, we ensure that the network not only meets the MLU optimization objectives but also enhances the burst resistance of different SD pairs in a fine-grained manner.